# Multipole Attention for Efficient Long Context Reasoning

### v1

@@ 5.2 Evaluation
Figure 4 presents evaluation on the LongBenchV2 dataset for the Deepseek-R1-Distil-Qwen-14B model. These results show how Multipole Attention is able to preserve the accuracy of the baseline, even with aggressive sparsity settings. Additionally, these results highlight how our method outperforms existing sparse attention methods, demonstrating how multipole attention approximation is able to retain closer accuracy to the baseline by leveraging the scores to the centroids. Table 1 also presents evaluation on LongBenchV2 for the Qwen3-8B model. We include baseline comparisons with QUEST Tang et al. (2024) and Squeezed Attention, demonstrating how Multipole Attention outperforms existing sparse attention methods at different token budgets. Note that we only report accuracy on the “Short” (<32K) split from the dataset, since the Qwen3-8B model achieves lower than 25% accuracy on the medium / long splits (which is worse than random since the questions are multiple choice with 4 answers), primarily due to issues with endless repetitions.
Table 2 provides evaluation for the Qwen3-8B model on GSM-Infinite. We include results for the single-level and hierarchical variants of our method and the Squeezed Attention baseline Hooper et al. (2024a). We report results for 1 and 2 operation difficulty levels (referring to the number of operations required to get to the final answer). Our method retains closer to baseline accuracy than sparse attention methods like Squeezed Attention. The advantages of Multipole Attention are particularly pronounced for aggressive sparsity regimes since there is a greater chance that sparse attention methods like Squeezed Attention will discard tokens which are important for retaining accuracy, whereas our approach always ensures that we retain at least an approximate representation for these tokens. Additionally, our hierarchical multipole algorithm attains higher accuracy for the same number of memory operations required relative to the one-level approach by reducing the overhead of the centroid lookup and centroid replacement operations.
@@ 5.3 System Benchmarking
We benchmarked our custom kernel implementations on both A6000 and A100 GPU platforms. Figure 5 shows the observed attention speedups on an A6000 GPU with our method, and kernel benchmarking evaluation on an A100 GPU is also provided in Appendix B. We observe up to 2.5, 4.1, and 4.5 speedups for batch sizes of 1, 4, and 16, respectively. We also show the breakdown of the portion of the runtime spent on our initial centroid comparison (“Centroid Lookup”), attention approximation using the centroids (“Centroid Replacement”), sparse Flash Decoding for exact attention with the important keys (“Sparse FlashDecoding”), as well as the median overhead from the fast cluster update (“Cluster Update”). For larger batch sizes, the centroid lookup and sparse FlashDecoding runtimes are substantial relative to the attention approximation kernel and cluster update, demonstrating the low overhead of our centroid replacement method relative to sparse attention, and how our fast cluster update enables quickly re-building an index over the generated tokens. Appendix C provides additional analysis of the clustering overheads during prefill as well as decode.
@@ 7 Limitations
One limitation of our method is that it focuses on speeding up generation, and it does not accelerate attention during the prefill phase. One potential avenue for future work would be extending our method to support accelerating prefill. Another limitation is that additional implementation effort would be required to support our method in existing LLM serving frameworks. Finally, we evaluated our method primarily on reasoning models, and broader evaluation across a range of long context inference tasks is also outside the scope of this work.
@@ Appendix A Extended Related Work: Only Loading Important KV Cache Entries
One previous approach for sparse attention has aimed to retain the full KV cache, but only load in the required KV cache entries for each decoding step. A common method that these approaches have used to identify important KV cache entries is clustering the keys. QUEST Tang et al. [2024] grouped consecutive KV cache entries and derived representative vectors for the keys within each group. Their approach performs retrieval by comparing the current query with the representative vectors in order to identify whether the group is important and needs to be loaded for exact attention computation. While this method grouped consecutive keys, positional proximity in the sequence is not necessarily indicative of the similarity of the corresponding key vectors. Squeezed Attention Hooper et al. [2024a] instead clustered keys based on semantic similarity, thereby allowing for precise identification of tokens which were likely to be high-scoring. Tactic Zhu et al. [2025] also clustered the keys based on semantic similarity and then performed distribution fitting to help predict important KV cache entries.
Multiple prior works have framed the query-key comparison as a retrieval problem in order to leverage vector search methods to accelerate attention. One such method, PQCache Zhang et al. [2024a], performed product quantization in order to perform vector search to select important keys, and then only computed exact attention using the selected keys. A2ATS He et al. [2025] performed query-aware vector quantization in order to allow for accurate retrieval of important KV cache entries. RetrievalAttention Liu et al. [2024a] constructs a vector index over the keys in order to facilitate retrieval of important keys, and offloads the full KV cache and the vector index to the CPU to perform the lookup.
@@
Another related prior work is Kang et al. [2023], which utilized a Fast Multipole Method (FMM)-inspired approach for accelerating the attention computation in Transformers. FMM Coifman et al. [1993] is a related technique for accelerating simulations with N-Body problems which approximates groups of point masses as a single point mass and performs coarser grained approximations for masses which are further away from the current point of interest. This approach is analogous to our algorithm, which approximates keys using progressively coarser-grained key centroids as we get further from the current query token. However, unlike Kang et al. [2023] which grouped KV cache entries based on positional proximity in the sequence, our method clusters keys based on semantic similarity, and therefore determines the granularity of approximation based on the importance of KV cache entries for the current query.
Appendix B A100 Kernel Benchmarking
@@
Appendix C Clustering Runtime
Tables 3 and 4 provide measured runtime for the fast cluster update during generation on A6000 and A100 GPUs, respectively. We find that the overhead of the fast cluster update is typically 3-4% of the runtime of the FlashDecoding baseline on an A6000 GPU, and 5% of the runtime of the FlashDecoding baseline on an A100 GPU. Table 5 also reports the clustering overhead during prefill, which is relatively low (between 13-15% on A6000/A100 GPUs).

### v2

@@ 5.2 Evaluation
Figure 4 presents evaluation on the LongBenchV2 dataset for the Deepseek-R1-Distil-Qwen-14B model. These results show how Multipole Attention is able to preserve the accuracy of the baseline, even with aggressive sparsity settings. Additionally, these results highlight how our method outperforms existing sparse attention methods, demonstrating how our multipole attention approximation is able to retain closer accuracy to the baseline by leveraging the scores to the centroids. Table 1 also presents evaluation on LongBenchV2 for the Qwen3-8B model. We include baseline comparisons with QUEST (tang2024quest) and Squeezed Attention, demonstrating how Multipole Attention outperforms existing sparse attention methods at different token budgets. Note that we only report accuracy on the “Short” (<32K) split from the dataset, since the Qwen3-8B model achieves lower than 25% accuracy on the medium / long splits (which is worse than random since the questions are multiple choice with 4 answers), primarily due to issues with endless repetitions. Appendix B provides additional analysis on LongBenchV1 (bai2023longbench) using a non-reasoning long context model, demonstrating how our method is generalizable for long context tasks outside of reasoning. Appendix C also provides ablations for the block size and number of cluster centroids hyperparameters.
Table 2 provides evaluation for the Qwen3-8B model on GSM-Infinite. We include results for the single-level and hierarchical variants of our method and the Squeezed Attention baseline (hooper2024squeezed). We report results for 1 and 2 operation difficulty levels (referring to the number of operations required to get to the final answer). Our method retains closer to baseline accuracy than sparse attention methods like Squeezed Attention. The advantages of Multipole Attention are particularly pronounced for aggressive sparsity regimes since there is a greater chance that sparse attention methods like Squeezed Attention will discard tokens which are important for retaining accuracy, whereas our approach always ensures that we retain at least an approximate representation for these tokens. Additionally, our hierarchical multipole algorithm attains higher accuracy for the same number of memory operations required relative to the one-level approach by reducing the overhead of the centroid lookup and centroid replacement operations.
@@ 5.3 System Benchmarking
We benchmarked our custom kernel implementations on both A6000 and A100 GPU platforms. Figure 5 shows the observed attention speedups on an A6000 GPU with our method, and kernel benchmarking evaluation on an A100 GPU is also provided in Appendix E. We observe up to 2.5, 4.1, and 4.5 speedups for batch sizes of 1, 4, and 16, respectively. We also show the breakdown of the portion of the runtime spent on our initial centroid comparison (“Centroid Lookup”), attention approximation using the centroids (“Centroid Replacement”), sparse Flash Decoding for exact attention with the important keys (“Sparse FlashDecoding”), as well as the median overhead from the fast cluster update (“Cluster Update”). For larger batch sizes, the centroid lookup and sparse FlashDecoding runtimes are substantial relative to the attention approximation kernel and cluster update, demonstrating the low overhead of our centroid replacement method relative to sparse attention, and how our fast cluster update enables quickly re-building an index over the generated tokens. Appendix D compares the attention speedups from Multipole Attention with Squeezed Attention, demonstrating that we achieve comparable speedups for the same memory footprint. Appendix F provides additional analysis of the clustering overheads during prefill as well as decode.
@@ 7 Limitations
One limitation of our method is that it focuses on speeding up generation, and it does not accelerate attention during the prefill phase. One potential avenue for future work would be extending our method to support accelerating prefill. Another limitation is that additional implementation effort would be required to support our method in existing LLM serving frameworks. A third limitation is that our method requires additional memory to store the centroids (although this overhead is small relative to the size of the full KV cache).
@@ Appendix A Extended Related Work: Only Loading Important KV Cache Entries
One previous approach for sparse attention has aimed to retain the full KV cache, but only load in the required KV cache entries for each decoding step. A common method that these approaches have used to identify important KV cache entries is clustering the keys. QUEST [tang2024quest] grouped consecutive KV cache entries and derived representative vectors for the keys within each group. Their approach performs retrieval by comparing the current query with the representative vectors in order to identify whether the group is important and needs to be loaded for exact attention computation. While this method grouped consecutive keys, positional proximity in the sequence is not necessarily indicative of the similarity of the corresponding key vectors. Squeezed Attention [hooper2024squeezed] instead clustered keys based on semantic similarity, thereby allowing for precise identification of tokens which were likely to be high-scoring. Tactic [zhu2025tactic] also clustered the keys based on semantic similarity and then performed distribution fitting to help predict important KV cache entries.
Multiple prior works have framed the query-key comparison as a retrieval problem in order to leverage vector search methods to accelerate attention. One such method, PQCache [zhang2024pqcache], performed product quantization in order to perform vector search to select important keys, and then only computed exact attention using the selected keys. A2ATS [he20252] performed query-aware vector quantization in order to allow for accurate retrieval of important KV cache entries. RetrievalAttention [liu2024retrievalattention] constructs a vector index over the keys in order to facilitate retrieval of important keys, and offloads the full KV cache and the vector index to the CPU to perform the lookup.
@@
Another related prior work is [kang2023fast], which utilized a Fast Multipole Method (FMM)-inspired approach for accelerating the attention computation in Transformers. FMM [coifman1993fast] is a related technique for accelerating simulations with N-Body problems which approximates groups of point masses as a single point mass and performs coarser grained approximations for masses which are further away from the current point of interest. This approach is analogous to our algorithm, which approximates keys using progressively coarser-grained key centroids as we get further from the current query token. However, unlike [kang2023fast] which grouped KV cache entries based on positional proximity in the sequence, our method clusters keys based on semantic similarity, and therefore determines the granularity of approximation based on the importance of KV cache entries for the current query.
Appendix B Additional Long Context Experiments
To demonstrate the broader applicability of our methodology on long context tasks other than reasoning, we evaluate our method on LongBenchV1 [bai2023longbench] using a non-reasoning long context model (Llama-3.1-8B-Instruct [llama31]), with results provided in Table 3. We also provide additional baseline comparisons with Squeezed Attention [hooper2024squeezed], DuoAttention [xiao2024duoattention] and TOVA [oren2024transformers]. We use the evaluation setup from the DuoAttention open-source code, which simulates decoding for the last 50 prompt tokens, and we use a temperature of 0. We configure Multipole Attention to use 95% sparsity and one centroid per 16 tokens, and then adjust the other methods to have equivalent KV cache budgets (88.75% sparsity when accounting for metadata). Our results demonstrate how our method substantially outperforms Squeezed Attention, DuoAttention, and TOVA on a range of long context evaluation tasks, and also how our method generalizes to long context tasks outside of reasoning.
| Config | Mem. | Single-Doc. QA | Multi-Doc. QA | Summarization | Few-shot Learning | Code | Avg. | |||||||||
| Ops. |
NQA |
Qspr |
MFQA |
HPQA |
2Wiki |
MSQ |
GRep |
QMSM |
MNews |
TREC |
TQA |
SSum |
RB |
LCC |
| Baseline | 1.00 | 30.1 | 45.3 | 55.5 | 56.0 | 44.8 | 30.5 | 35.2 | 25.3 | 27.3 | 72.5 | 91.2 | 43.2 | 49.0 | 52.5 | 47.0 |
| DuoAttn | 0.1125 | 13.2 | 16.8 | 21.3 | 29.1 | 17.4 | 8.2 | 20.5 | 17.3 | 20.1 | 32.0 | 52.4 | 36.8 | 47.0 | 49.0 | 27.2 |
| TOVA | 0.1125 | 24.1 | 14.0 | 23.2 | 41.5 | 21.1 | 15.1 | 25.5 | 19.9 | 20.9 | 35.0 | 86.1 | 39.9 | 49.1 | 48.6 | 33.1 |
| SqAttn | 0.1125 | 29.3 | 24.2 | 38.6 | 50.1 | 35.2 | 27.0 | 31.5 | 24.5 | 18.9 | 55.0 | 81.6 | 41.5 | 49.9 | 47.4 | 39.6 |
| MpAttn | 0.1125 | 31.6 | 32.9 | 42.0 | 54.3 | 31.3 | 30.3 | 32.8 | 24.7 | 22.9 | 67.5 | 86.7 | 42.1 | 46.8 | 48.7 | 42.5 |
Appendix C Hyperparameter Ablations
Tables 4 and 5 provide ablations for the number of centroids as well as block size, respectively. We report LongBenchV2 accuracy for the Qwen3-8B model on the “short” (<32K context length) split with a token budget of 128. We find that if the number of KV tokens per centroid is increased or if the block size is decreased, the accuracy is degraded. However, increasing these parameters leads to additional inference costs. We therefore used 1 centroid per 16 tokens and a block size of 8K to retain model accuracy without introducing substantial memory and latency overheads.
| Baseline | Ratio=1/128 | Ratio=1/64 | Ratio=1/32 | Ratio=1/16 | Ratio=1/8 |
| 41.5 | 29.3 | 35.4 | 38.5 | 38.1 | 37.0 |
| Baseline | ||||
| 41.5 | 35.0 | 38.0 | 38.1 | 36.9 |
Appendix D Runtime Comparison with Squeezed Attention
| Method | Batch Size = 1 | Batch Size = 4 | Batch Size = 16 |
| Squeezed Attention | 2.8 | 3.3 | 3.6 |
| Multipole Attention | 2.4 | 3.3 | 3.7 |
In Table 6, we compare the latency for Multipole Attention versus Squeezed Attention [hooper2024squeezed] for the same memory footprint (using one centroid per 16 tokens for Multipole Attention and one centroid per 8 tokens for Squeezed Attention, as outlined in Section 5.1). Note that the Squeezed Attention baseline we compare with in our experiments also contains the clustering improvements from Multipole Attention, which are required to allow it to be applied for tasks where the prefill is only known at runtime and during the generation process to accelerate attention to newly generated tokens (as the baseline Squeezed Attention method cannot be run online). We report the speedups for Squeezed Attention and Multipole Attention on an A6000 GPU with batch sizes of 1/4/16 (assuming 90% sparsity). These results outline how Multipole Attention provides similar latency for the same sparsity level as Squeezed Attention, while providing substantial accuracy improvements (as highlighted in Section 5.2).
Appendix E A100 Kernel Benchmarking
@@
Appendix F Clustering Runtime
Tables 7 and 8 provide measured runtime for the fast cluster update during generation on A6000 and A100 GPUs, respectively. We find that the overhead of the fast cluster update is typically 3-4% of the runtime of the FlashDecoding baseline on an A6000 GPU, and 5% of the runtime of the FlashDecoding baseline on an A100 GPU. Table 9 also reports the clustering overhead during prefill, which is relatively low (between 13-15% on A6000/A100 GPUs).